Check the alarm on U2000, no R_LOS or TU_AIS alarm on the RTN in site B
Check the configuration for the E1 service, it was ok
Software loop back on the RTN in site B
, check the E1 status from BSC. The E1 status was degraded.
Check the MW link performance, it was ok.
Check the input and out power of the optical board, which connected from OSN to RTN and from OSN to BSC, it was on the board range.
Software loop back from OSN to the BSC direction on the RTN optical port.
Check the E1 status on the BSC, it was Available.
Hardware loop back from OSN to the BSC direction on the RTN optical port.
Check the E1 status on the BSC, it was degraded.
Keep the hardware loop and check the power of the optical port, it was normal.
Keep the hardware loop and Shutdown the laser of the optical port, then check the input output power,
we found that the port outpower is -60dbm, but the Inpower is -10dbm,